# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 5.2.1 HEK293 data
RESET Ensemble 27.5 minutes
AdaKO GAM 1.06 days
AdaKO EM 33.71 days
AdaKO LR 3.35 days
AdaKO RF 836.03 days
Table 2: The average runtimes for each method at 1% FDR on the HEK293 dataset. The average runtimes reported for each Adaptive Knockoffs method is estimated (as described in Section S.13), while we computed the actual average runtime for RESET Ensemble. The average is taken over 5 randomly constructed decoy databases. We used 20 cores for RESET Ensemble.
@@
Finally, using the method described in Supplementary Section S.13, we estimated the time it would take for each method of Adaptive Knockoffs to report discoveries at the 1% FDR level (which is the typical FDR threshold in the peptide detection context). Table 2 shows the average estimated times taken over 5 distinct target-decoy databases, with decoys varying each time by randomly shuffling the target peptides. It is clear that running any implementation of AdaKO is generally too expensive, with the worst being AdaKO RF at an estimate of 836 days, and the default method, AdaKO GAM, taking an estimate of 1.06 days. In contrast, we were able to successfully run RESET Ensemble with FDR control in 27.5 minutes.
5.2.2 PRIDE data
Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context. Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the twenty PRIDE-20 datasets that we recently used in [18]. In the following results, we only considered searching the spectrum files once against a single target-decoy database, since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs. Data preparation, search settings and the detailed definition of the triples $(L,W,\mathbf{x})$ are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.555The side information here slightly differs from above due to different search settings.
@@ 5.2.2 PRIDE data
Figure 5: Comparison of RESET Ensemble and AdaKO GAM. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
@@
Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot). In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets. Clearly, RESET Ensemble is overall more powerful than AdaKO GAM and TDC. Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is reporting fewer discoveries than the generic TDC.
In a second comparison, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs. In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF. In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced more discoveries. In one dataset, AdaKO RF discovered 1217 peptides while RESET Ensemble only discovered 1104 peptides, about a 10% increase over RESET Ensemble.
We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13. Reassuringly, our estimated times were less than the actual times taken by each method to complete, thus validating the conservative nature of our estimated times in the previous section. Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets. In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 10%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete. Figure 6C shows that in the worst case, RESET takes roughly 1 minute. The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analyzed jointly, as in our HEK293 analysis in the previous section.
Figure 6: Comparison of RESET Ensemble and AdaKO RF. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The left panel shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO RF (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot). In the centre panel, we compare the estimated and actual runtimes for AdaKO RF. Lastly, in the right panel, we compare the actual runtimes of RESET Ensemble and AdaKO RF (in log10 scale).
5.2.3 FDP control
Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and all twenty PRIDE-20 spectrum files from the previous sections. As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides. Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases. We used the coinflip version of FDP-SD [40], which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm S2). Our results using RESET Ensemble and FDP-SD are given in Table 3 and Figure 7.
Confidence $1-\gamma$
Method 0.5 0.8 0.9
RESET Ensemble 82987 82729 82558
FDP-SD 75977 75464 75254
Table 3: The average number of discoveries at 1% FDP threshold at varying confidence parameters using the HEK293 data. The average number of discoveries reported for RESET Ensemble and FDP-SD on the combined HEK293 spectrum files at different confidence values $0.91-\gamma=0.5,0.8,$ 0.9. The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful, with an approximately 9–10% increase in the average number of discoveries over FDP-SD at all confidence levels. Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets it fails to make any discoveries at high confidence levels. Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417. When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in the Discussion. Regardless, more often than not RESET Ensemble produces more discoveries, and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
Figure 7: Comparison of RESET Ensemble and FDP-SD. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the twenty PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
The results above consider a canonical ‘narrow’ search. Specifically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the observed mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount. Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30]. Such a search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs). We provide further detail in Supplementary Section S.14. In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs. Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context (Figure 8).
Figure 8: Comparison of RESET ensemble and FDP-SD using the recently popular ‘open’ search. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. We implement an ‘open’ search for each PRIDE-20 dataset (see Section S.14). The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
@@ 6 Experiments: p-value setting
6.1 Numerical experiments
In this section, we consider the p-value setting, comparing RESET with the FDR controlling procedures we reviewed in Section 3.1 on a set of simulations introduced by Lei and Fithian [33]. As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. Details of each method used are given in Supplementary Sections S.6S.10.
@@ 6.1.1 Simulation 5: p-value based testing with geometric side information
In this section, we look at Simulation 1 of Lei and Fithian [33]. For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis. A total of $m=2500$ hypotheses are used. Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$. The false null hypotheses correspond to certain regions in the square. They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure 9). We used 100 independent runs, where in each run we randomly draw the 2500 p-values as described.
Figure 9: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same picture is given in [33].
Figure 10: RESET vs. other methods using two-dimensional side information with p-values. The first row plots the power and the second row plots the estimated FDR for each FDR threshold ranging from 1% to 20%. Each color and shape combination corresponds to a unique method. The three columns correspond to the hypotheses’ relationship to the side information according to Figure 9. The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
Figure 10 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1. We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal to the best method in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM. However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels. At 1% FDR, all RESET methods have negligible power, although this is consistent with many of the other methods. Interestingly, many of the finite-FDR controlling procedures such as the various RESET and AdaPT methods generally have better power than the asymptotic approaches, ZAP and AdaFDR, suggesting perhaps a misspecification of the fitted model used by these approaches. Figure 10 also shows the estimated FDR of each method. We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold. This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control. Consistently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
@@ 6.1.2 Simulation 6: p-value based testing with 100-dimensional side information
In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d. draws from $U[0,1]$. This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses. Each p-value $p_{i}$ is drawn i.i.d. from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution. Specifically, the beta mixture model is given by:
where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$. The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
where $\theta=(3,3,0,\dots,0)\in\mathbb{R}^{100}$, $\beta=(2,2,0,\dots,0)\in\mathbb{R}^{100}$ and $\theta_{0}$ is chosen to satisfy $\frac{1}{m}\sum_{i}\pi(\mathbf{x}_{i})=0.3$. Clearly, only the first two values in $\mathbf{x}_{i}\in[0,1]^{100}$ contribute to the beta mixture model, as intended. We used 100 independent runs, where in each we drew the 2000 p-values as described.
Figure 11: RESET vs. other methods using 100-dimensional side information with p-values. The left panel plots the power and the right panel plots the estimated FDR as a function of the FDR thresholds: 1%, 5%, 10% and 20%. Each color and shape combination corresponds to a unique method. The horizontal segments in the right panel mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
@@
Examining Figure 11, which graphically summarizes this analysis, it appears that RESET Ensemble, RESET NN, and AdaPT GLMnet deliver the most discoveries at the 5%, 10%, and 20% FDR levels. At 1% FDR, only AdaPTg GLMnet appears to have non-negligible power, albeit a fairly low number of discoveries. Finally, we point out that in this simulation, the model specification of AdaPT/AdaPTg perfectly matches the true model. In other words, the fitted mixture model used by AdaPT/AdaPTg coincides with Equation (6). Given this, we find it reassuring that RESET Ensemble is still comparable at 5% and higher thresholds.

### v3

@@ 5.2.1 Runtime comparison
RESET Ensemble 17.7 minutes
AdaKO GAM 1.07 days
AdaKO EM 34.17 days
AdaKO LR 3.50 days
AdaKO RF 846.33 days
@@
We estimated the time it would take for each method of Adaptive Knockoffs to report discoveries at the 1% FDR level (see Section C.14). Table 2 shows the estimated times averaged over 5 different constructions of the decoy database by randomly shuffling the target peptides. All the implementations of AdaKO are not practical for analyzing this data, with the worst being AdaKO RF at an estimated 846.33 days, and the default method, AdaKO GAM, taking an estimate of 1.07 days. In contrast, we were able to run RESET Ensemble in 17.7 minutes on average using 20 cores (Section C.13).
5.2.2 Power comparison
General applicability aside, we resorted to using small datasets to evaluate the power of Adaptive Knockoffs in the peptide detection context. Specifically, we focused on the thirteen smallest datasets from PRIDE-20, a collection of twenty datasets that we recently used (Freestone et al. (2024)). For each of the thirteen spectrum files in PRIDE-20, we performed two types of searches: a narrow and an open search. Briefly, in a narrow search, spectra can only match database peptides that are within a small tolerance of the machine-reported mass of the molecules that generated those spectra. On the other hand, in open searches the same tolerance is several orders of magnitude larger. Details regarding PRIDE-20 and the search settings are given in Section D.2.
@@ 5.2.2 Power comparison
Figure 2: Comparison of RESET Ensemble to AdaKO GAM and TDC. For each of the PRIDE-20 datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplots) and TDC (right boxplots) applied to thirteen PRIDE-20 datasets and two different search modes (there are total of 13 points in each boxplot).
@@
Figure 2 shows the average number of discoveries reported using RESET Ensemble divided by the number of discoveries reported by AdaKO with its default GAM filter (left boxplots) and TDC (right boxplots), using the two different search settings. We averaged RESET Ensemble over 10 applications by varying the internal seed which randomly splits the decoys into their training and estimating sets. Clearly, RESET Ensemble is overall more powerful than AdaKO GAM and TDC. Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is at times reporting fewer discoveries than the generic TDC.
In Section E.2, we compare RESET Ensemble with AdaKO RF since we found AdaKO RF to be the most powerful variant of Adaptive Knockoffs in this setting. The results show that RESET Ensemble is comparable to AdaKO RF in terms of power, with the latter still being generally too slow to be practical.
@@ 6 Experiments: p-value setting
6.1 Assumption guaranteeing finite-sample FDR control
Recall that RESET and Assumption 3 that guarantees its finite-sample error control are formulated in terms of the labels and winning score $(L,W)$. To apply RESET in the p-value setting, we first convert each $p_{i}$ into a pair $(L_{i},W_{i})$ as in Equation (4). Similarly, instead of Assumption 3 we adopt the following assumption formulated by Chao and Fithian (2021) who also showed it implies the former assumption (the next lemma).
Assumption 4.
Let $N$ be the indices of the true null hypotheses. Conditioned on the side information $\mathbf{x}$, and the false null p-values $(p_{i})_{i\not\in N}$, the true null p-values $(p_{i})_{i\in N}$ are independent and have a non-decreasing density.
Lemma 2 (Chao and Fithian).
Assumption 4 implies Assumption 3 with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$, where $[0,a)$ and $(b_{1},b_{2}]$ are the regions corresponding to the positive and negative labels as described in Equation (4).
The above lemma is phrased differently by Chao and Fithian (2021) and is presented as Lemma A.1 in their paper, but it is materially the same. Most of the time, we will consider symmetric cutoff regions $[0,1/2)$ and $(1/2,1]$. For completeness, we provide a proof of this lemma in Section B.2 using our notation.
6.2 Simulated data
In this section, we consider the p-value setting, comparing RESET with the procedures we introduced in Section 3.1 on a set of simulations introduced by Lei and Fithian (2018). Each setup, detailed through Sections C.7C.11, is designed to satisfy Assumption 4 and hence Assumption 3, which is required for our theoretical guarantees.
@@ 6.2.1 Simulation 5: p-value based testing with geometric side information
In this section, we look at Simulation 1 of Lei and Fithian (2018). For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis. A total of $m=2500$ hypotheses are used. Each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$. The false null hypotheses correspond to certain regions in the square. They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure C12). We generated and analyzed 100 independent datasets, where in each we randomly draw the 2500 p-values as described.
Figure 3: RESET vs. other methods using p-values with two-dimensional side information. The power (1st row) and the estimated FDR (2nd row) are plotted against FDR thresholds ranging from 1% to 20%. Each color and shape combination corresponds to a unique method. The three columns correspond to the hypotheses’ relationship to the side information according to Figure C12. The black horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
Figure 3 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1. We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal to the best method in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM. However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels. At 1% FDR, all RESET methods have negligible power, although this is consistent with many of the other methods. Interestingly, many of the finite-FDR controlling procedures such as the various RESET and AdaPT methods generally have better power than the asymptotic approaches, ZAP and AdaFDR, suggesting perhaps a misspecification of the fitted model used by these approaches. Figure 3 also shows the estimated FDR of each method. We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold. This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDX control. Consistently, Chao and Fithian (2021) reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
@@ 6.2.2 Simulation 6: p-value based testing with 100-dimensional side information
We next adopted Simulation 2 of Lei and Fithian (2018). The side information for $m=2000$ hypotheses consists of 100 i.i.d. draws from $U[0,1]$ in which only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses. Details are given in Section C.12. We conducted 100 independent runs, where in each run we randomly draw the $m=2000$ p-values.
Figure 4: RESET vs. other methods using p-values with 100-dimensional side information. The power (left panel) and the estimated FDR (right panel) are plotted as a function of the FDR thresholds: 1%, 5%, 10% and 20%. Each color and shape combination corresponds to a unique method. The black horizontal segments in the right panel mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
@@
Examining Figure 4, which graphically summarizes this analysis, it appears that RESET Ensemble, RESET NN, and AdaPT GLMnet deliver the most discoveries at the 5%, 10%, and 20% FDR levels. At 1% FDR, only AdaPTg GLMnet appears to have non-negligible power, albeit a fairly low number of discoveries. Finally, we point out that in this simulation, the model specification of AdaPT/AdaPTg perfectly matches the underlying data generating model: the fitted mixture model used by AdaPT/AdaPTg coincides with Equation (7) of Section C.12. Given this, we find it reassuring that RESET Ensemble is still comparable at 5% and higher thresholds.
@@
6.3 Robustness to dependent data
In reality, real datasets often exhibit varying degrees of dependency, therefore failing to satisfy Assumptions 3 and 4. RESET Ensemble is equipped with a heuristic to account for such data (Section C.2.4). The following simulations adapted from Zhang et al. (2019) empirically demonstrate that RESET Ensemble exhibits some robustness to both weak and strong forms of local dependency.
6.3.1 Simulation 7: Weak dependent p-values
This simulation is based on Data 4 of Zhang et al. (2019). We generated $m=20K$ hypotheses with i.i.d side information drawn from $\mathbf{x}_{i}\sim U[0,1]$. Each hypothesis was declared a false null with probability $\pi_{i}:=\pi(\mathbf{x}_{i})$ which is dependent on the drawn side-information. The function $\pi$ is a convex combination of (truncated) normal densities and an exponential function (Section C.12.2). For every 10 consecutive hypotheses, starting with $i=1,\dots,10$, we drew $z_{i}$ jointly from a $\mathcal{N}(0,\Sigma)$ where $\Sigma$ is a covariance matrix with upper triangular matrix given by $\Sigma_{jj}=1$, $\Sigma_{jk}=0.25$ for $1\leq j<k\leq 5$, $\Sigma_{jk}=-0.25$ for $1\leq j\leq 5<k\leq 10$ and $\Sigma_{jk}=0$ for all other entries in the upper triangle. Hypotheses which are false nulls had their z-scores shifted by 2. A one-sided p-value $p_{i}=1-\Phi(z_{i})$ is then obtained. Notably, the covariance matrix $\Sigma$ induces a local weak dependency among the p-values, effecting both true and false null hypotheses.
6.3.2 Simulation 8: Strong dependent p-values
This simulation is based on Data 5 of Zhang et al. (2019). We used $m=20K$ hypotheses with side information $\mathbf{x}_{i}$ and false null probability $\pi_{i}$ defined as in Simulation 7. For every 5 consecutive hypotheses, starting with $i=1,\dots,5$, $z_{i}=z$ where $z\sim\mathcal{N}(0,1)$, creating 5 identical z-scores. False null hypotheses are shifted by 2 and one-sided p-values $p_{i}=1-\Phi(z_{i})$ were obtained. In this case, the data exhibits a local strong dependency effecting both true and false null p-values.
6.3.3 Results
Figure 5 shows the estimated FDR calculated from 100 runs of the above simulations. In the weakly dependent case (A), the majority of the methods seem to empirically control the FDR, with the exception of AdaFDR and RESET GAM which appears to marginally exceed the 5% threshold. In the strongly dependent data (B), we see several more methods exceeding their respective thresholds, albeit relatively mildly. RESET Ensemble seems to be robust and appears to approximately control the FDR in these experiments.
Figure 5: Estimated FDR using RESET vs. other methods on dependent p-values. The estimated FDR is plotted as a function of the FDR thresholds: 1%, 5%, 10% and 20% for weak (A) and strong dependence (B). Each color and shape combination corresponds to a unique method. The black horizontal segments mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.